For project A, the cash flow effect from the change in net working capital is expected to be 100 dollars at time 2 and the level of net working capital is expected to be 1,800 dollars at time 1. What is the level of current assets for project A expected to be at time 2 if the level of current liabilities for project A is expected to be 1,600 dollars at time 2?
Answer #1

Net Working Capital = Current Assets - Current Liabilities

Change in net working capital = Current net working capital - Previous net working capital

Change in net working capital at Time-2 = 100

Net working capital at Time-1 = 1,800

Thus,

Net working capital at Time-2 = 100+1,800 = 1,900

Expected current liabilities at Time-2 = 1,600

Therefore,

\textup{Expected Current Assets for Project A at Time 2 = 1,900+1,600 = {\color{Green} \textbf{\$3,500}}}

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.